### v1

# MonoSOWA: Scalable monocular 3D Object detector Without human Annotations

Abstract
Detecting the three-dimensional position and orientation of objects using a single RGB camera is a foundational task in computer vision with many important applications. Traditionally, 3D object detection methods are trained in a fully-supervised setup, requiring vast amounts of human annotations, which are laborious, costly, and do not scale well with the ever-increasing amounts of data being captured.
In this paper, we present the first method to train 3D object detectors for monocular RGB cameras without domain-specific human annotations, thus making orders of magnitude more data available for training. Thanks to newly proposed Canonical Object Space, the method can not only exploit data across a variety of datasets and camera setups to train a single 3D detector, but unlike previous work it also works out of the box in previously unseen camera setups. All this is crucial for practical applications, where the data and cameras are extremely heterogeneous.
The method is evaluated on two standard autonomous driving datasets, where it outperforms previous works, which, unlike our method, still rely on 2D human annotations. The source code and the model is going to be published soon.
(a) Traditional pipelines require human labelling and produce new model for every different camera setup
(b) Our method does not need human labelling and combines multiple camera setups into a single model operating in canonical object space
Figure 1: Traditional pipelines (a) vs. the proposed pipeline (b) for monocular 3D object detection.
I Introduction
Monocular 3D object detection is a key component of many important systems, ranging from robotics to autonomous cars. 3D object detection methods are typically trained in a fully-supervised setup, requiring vast amounts of human annotations, which are laborious, costly, and do not scale well with ever-increasing amounts of data that ideally should be used for training. To make matters even worse, especially in autonomous driving, new cameras and new camera setups emerge constantly, which typically implies the need to collect, label and train on a new dataset as camera parameters such as focal length change.
In this paper, we address both of these challenges – costly human labels and changing camera setup – by introducing a new method to train a 3D rigid object detector in a Canonical Object Space without domain-specific human labels. Our method allows us to exploit large amounts of training data readily available because all we need as input for training are unlabelled driving video sequences. Furthermore, our method does not require data from additional sensors such as LiDAR. It makes our method extremely applicable in practice because as such our method can directly exploit data collected by a majority of currently manufactured vehicles, which typically only have a single camera. This is in contrast with some methods which require LiDAR (and human annotations) [1, 2], because for them specialised vehicles are needed to collect the data, and this severely limits the variety of geographies, environments, and traffic situations being captured.
Drawing inspiration from monocular depth literature [3], the proposed Canonical Object Space (COS) allows us to aggregate and learn from data coming from different cameras, which is a crucial pre-condition to being able to effectively harvest otherwise very heterogeneous video sequences. Moreover, unlike previous methods such as VSRD [4], it also allows for the deployment of our 3D object detector in new unseen camera setup without needing to train or fine-tune for the new setup. This is again extremely useful in practice, because new camera sensors and setups are developed all the time, but it is unfeasible to collect data, label them by human annotators, and then train a new model for each new camera type.
To summarize, we make the following contributions:
  1. 1.
    We propose a novel auto-labelling pipeline which exploits temporal consistency in video sequences to automatically generate 3D position, size, and orientation labels of rigid objects (cars) from a monocular camera.
  2. 2.
    To the best of our knowledge, our method is the first method to train a monocular 3D object detector for autonomous driving which does not require neither 2D nor 3D human annotations; we demonstrate to the community that such an approach is feasible and is worth pursuing further.
  3. 3.
    We propose novel Canonical Object Space (COS), which allows us to a) aggregate data from different camera setups in training, and b) at inference time, use an already trained model in the previously unseen camera setup, i.e. without any training or fine-tuning for the new setup.
  4. 4.
    Our auto-labelling method is significantly faster than previous work, allowing us to process significantly more training data than the current state-of-the-art method.
II Related work
Monocular Depth Estimation
Monocular depth estimation has shown significant improvements in recent years. MiDAS [5] employs affine-invariant losses to enable training on multiple datasets. ZoeDepth [6] combines training on multiple datasets with relative losses and further fine-tunes on the target metric dataset. Depth Anything [7] focuses on the affine-invariant depth estimation using self-supervised pre-training. For metric depth estimation, which is relevant for object detection, it fine-tunes on the target dataset such as KITTI [8]. Metric3D [9, 3] focuses on metric depth estimation as it uses a Canonical Camera Transformation Module to further improve zero-shot performance, together with DINOv2 [10] or ConvNext [11] as the backbone and DPT [12] as the predictor. CamConv [13] models cameras inside the network, but as shown in [3], this leads to poor results.
Fully-supervised Monocular 3D Object Detection
Early methods [14, 15, 16] were based on lifting the image into pseudo-LiDAR and then detecting objects in the generated point cloud. SMOKE [17], on the other hand, directly predicts 3D bounding boxes from images. MonoDTR [18] and MonoDETR [19] use an end-to-end depth-aware/guided transformer [20] architecture to directly predict 3D bounding boxes. SSD-MonoDETR [21] introduces scale awareness into the model. MonoATT [22] variably assigns tokens to areas of higher importance. MonoFlex [23] uses an ensemble of direct depth and keypoint-based predictions and decouples the detection of truncated and untruncated vehicles. MonoCD [24] enhances the ensemble with a complementary depth predictor. However, all the aforementioned methods require 3D human annotations during their training.
Weakly-supervised 3D Object Detection
Most previous works on weakly-supervised 3D detection have focused on training 3D detectors for LiDAR. VS3D [25] pioneered in this area by using point density in LiDAR scans to identify objects, and detections are further refined by a student-teacher network, where the teacher is trained on images only. Zakharov et al. [2] employs a 2D off-the-shelf detector accompanied by a novel Signed Distance Fields (SDF)-renderer within the DeepSDF framework [26] and Normalized Object Coordinate Space (NOCS). The method is trained on a synthetic dataset and further fine-tuned on real datasets.
WS3D [27, 28] employs Bird’s Eye View (BEV) human-centered clicks as weak supervision to train a network for detecting vehicles, arguing that these are much easier to obtain than 3D bounding boxes. Furthermore, they use a small subset of human ground-truth annotations to train a second network to predict the spatial dimensions of the detected vehicles. McCraith et al. [29] employ direct fitting of generic templates into LiDAR scans segmented by Mask-RCNN [30] and share the information between frames by requiring consistency between frames while simultaneously discounting outliers. TCC-det [31] uses direct fitting of generic templates on aggregated points corresponding to each instance. During the training of the 3D detector, TCC-det appends two additional losses for finer alignment. VG-W3D [32] encodes the image and LiDAR scan separately and then enforces alignment between 2D and 3D on multiple feature levels to predict precise 3D bounding boxes.
Figure 2: Our auto-labelling pipeline creates 3D pseudo-groundtruth using only images, camera parameters and IMU data as inputs.
In contrast to the aforementioned methods, VSRD [4] focuses on 3D detection in camera, and does not require LiDAR scans. However, it uses 2D human ground-truth masks for each instance, representing them as a surface in Signed Distance Fields (SDF). Furthermore, the SDF are rendered into masks through their proposed volumetric-aware silhouette rendering, enabling end-to-end training. Each instance is optimized over multiple frames. While VSRD shows promising results in both weakly-supervised and semi-supervised training, it requires 2D instance masks, and rendering of each frame takes approximately 15 minutes. In contrast, our method does not require any human annotations at all and is approximately 100 times faster.
III Method
Traditionally, deep networks that predict 3D bounding boxes of vehicles (and other agents) from a single image are trained using 3D annotations, which are manually created by human annotators. They typically rely on both camera image and LiDAR data for correct distance estimate in creating their labels. However, this process is very laborious and costly, limiting the amount of training data available to train such methods. Moreover, it requires capturing the scene using both a camera and a LiDAR sensor, which is a significant limitation since the majority of production vehicles these days do not have LiDAR on board – as a result, data collection is only limited to vehicles dedicated to such a purpose, and this again limits the variety and diversity of data being used to train these methods.
Our method overcomes these limitations by introducing a novel approach to train a 3D object detector using only images (and vehicle motion data) as inputs, without requiring human labels or LiDAR in the training process.
III-A Auto-labelling pipeline
The process of automatic label creation assumes that images are captured in a sequence, which is a natural way to capture data in driving scenarios. It also assumes that the camera setup is known (both intrinsics and extrinsics) for each sequence. However, the camera setup might be and actually often is different (see section III-B). Lastly, it assumes approximate data about ego-vehicle’s motion is available (e.g. from GPS or an Inertial Measurement Unit – IMU), which again is commonplace.
Because labelling a 7-DOF pose of objects jointly is a very challenging problem, we instead infer individual pose parameters separately in a sequence of steps.
Detecting objects in pseudo-LiDAR
The auto-labelling process begins by inferring a metric depth map $D\in\mathbb{R}^{h\times w}$ for each camera image $I\in\mathbb{R}^{3\times h\times w}$, where $h$ and $w$ denotes image height and width, using an off-the-shelf monocular depth estimator. We opted to use Metric3D [3], as it is zero-shot, it does not limit our method in terms of data distribution and has shown great results for generalization. The choice of a particular monocular metric depth estimator is, however, not crucial. For each pixel and its corresponding depth, a 3D point cloud is generated as
$X_{u,v}$ $=\frac{D_{u,v}\cdot(u-c_{x})}{f_{x}}$
$Y_{u,v}$ $=\frac{D_{u,v}\cdot(v-c_{y})}{f_{y}}$
$Z_{u,v}$ $=D_{u,v}$ (1)
where $u\in[0,h]$ and $v\in[0,w]$, $c$ is the principal point and $f$ is thefocal length. We denote the pseudo lidar point cloud as $P_{i}\in\mathbb{R}^{3\times hw}$, where $i$ stands for the frame index.
Figure 3: Combining instance segmentation in image (top) with monocular metric depth estimation (middle) enables us to produce pseudo-LiDAR point cloud for each object.
Next, inspired by [29], we employ an off-the-shelf 2D object image detector to detect the objects we are interested in (vehicles) and their instance segmentation masks in each camera image. Given an instance segmentation mask $M_{i,j}\in\mathbb{R}^{h\times w}$ of an object $j$ in a frame $i$, we find object point cloud $P_{i,j}$ in the point cloud as
$P_{i,j}=\big{\{}p\in P_{i}\,|\,\mathcal{T}(p)\in M_{i,j}\big{\}}$ (2)
where $\mathcal{T}(p)$ denotes the projection of the 3D point $p$ into image coordinates (using camera calibration, which is known).
The $P_{i,j}$ tends to contain outliers, as either the masks or the metric depth estimation is not perfect. To remove the outliers, an ensemble of the following methods is employed: HBOS [33], Z-scores, statistical outlier rejection in Open3D [34], HDBSCAN [35] and DBSCAN [36]. If at least two of these methods classify a point as an outlier, it is removed. To compensate for the effect of the ego-vehicle movement, we transform all $P_{i,j}$ into the reference frame coordinate space, given the poses from IMU data.
Temporal consistency
In order to exploit temporal consistency across multiple frames of the sequence, we need to first establish correspondences of individual object instances between frames. Because we have inferred approximate 3D point cloud of each object instance $P_{i,j}$ in each frame, we track each object in 3D world coordinate system.
For each object $j$ in frame $i$, we approximate its location $L_{i,j}$ as the median of its point cloud $P_{i,j}$. The tracking is initialized in $n$ frames before the reference (current) frame, and for each frame, it matches the instances based on locations in the current frame $L_{i,j}$ and the predicted future locations from the last frame into the current frame $\hat{L}_{i,j}$. For the location prediction $\hat{L}_{i,j}$, a simple physics-aware motion model is used
$\hat{L}_{i,j}$ $=L_{i-1,j}+\frac{1}{3}\sum_{k=1}^{3}L_{i-k,j}-L_{i-k-1,j}$ (3)
To match two object instances between frames, it is required that the instances are both their nearest neighbours and that the distance between them is lower than a set threshold $T_{\text{dist}}$. Otherwise, the instance is considered as a new object. After this step, each object instance is represented by extracted points $P_{i,j}$ and locations $L_{i,j}$ in each frame the instance is present. The tracking sequentially processes all the frames up to $n$ after the reference frame.
Movement Classification
The next step is to classify instances as either stationary or moving, as different principles of temporal consistency exploitation are applied to those classes. Note that all object instances actually appear moving – they change location in relative terms, as the ego-vehicle itself is driving while data is captured.
In order to classify a vehicle as stationary or moving, we propose the following novel classification procedure. For each instance, all locations $L_{m...n,j}$ are taken and differences between all adjacent locations are computed, where $m$ is the first frame, where the instance is present and $n$ is the last. From those differences, the standard deviation $\sigma$ is calculated as
$\sigma_{j}$ $=\frac{1}{\sqrt{2}}\sqrt{\frac{1}{n-m}}\sum_{i=m+1}^{n}(\mu_{j}-D_{i,j})^{2}$ (4)
$\mu_{j}$ $=\frac{1}{n-m}\sum_{i=m+1}^{n}D_{i,j}\qquad D_{i,j}=L_{i,j}-L_{i-1,j}$ (5)
Furthermore, the actual distance $d$ and expected distance $\hat{d}$ caused by the $\sigma$ is calculated as
$d_{j}=\left\lVert L_{m,j}-L_{n,j}\right\rVert^{2}_{2}\qquad\hat{d}_{j}=(n-m)\cdot\left\lVert\sigma_{j}\right\rVert^{2}_{2}$ (6)
The $z$- and $p$-scores are calculated to know if the difference between the $d_{\text{real}}$ and $d_{\text{exp}}$ is statistically significant or not
$z_{j}=\frac{d_{j}}{\hat{d}_{j}}\qquad p_{j}=1-\Phi(z_{j})$ (7)
To classify the instance as moving, the $p$-score is required to be lower than the threshold $T_{m}$ and simultaneously, the $d$ distance must be higher than a given threshold.
As all $P_{i,j}$ are in the reference frame, we can aggregate points for stationary instances by simply concatenating the points
$A_{j}=\bigcup_{i=m}^{n}P_{i,j}$ (8)
Aggregating points for moving instances poses a significant challenge, as the $P_{i,j}$ differs significantly between frames. For that reason, for moving instances, only the trajectory and $P_{i,j}$ are employed.
Refer to caption Refer to caption
a) b)
Refer to caption Refer to caption
c) d)
Figure 4: Temporal aggregation examples. The point cloud of a significantly occluded vehicle (a) is recovered in subsequent frames (b) where the vehicle is no longer occluded. Distant car with poor depth prediction (c) is refined as the ego-vehicle drives closer (d). White points is pseudo-LiDAR of each frame, colourful points are the aggregated points. Human ground truth overlaid as green boxes for reference purposes.
Orientation and size estimation
Finding oriented 3D bounding boxes is a 7-degrees of freedom (DOF) problem. Instead of solving it directly, our method divides it into three disjunctive problems: orientation, location and dimensions.
Orientation estimation is different for moving and stationary instances. For moving instances, the yaw of the object is estimated by directly computing the angle between adjacent locations as
$\theta_{i,j}=\operatorname{atan2}\left(\frac{L_{i,j}(z)-L_{i-1,j}(z)}{L_{i,j}(x)-L_{i-1,j}(x)}\right)$ (9)
It is worth noting again that all locations are transformed into the reference coordinate space. To make it more robust, yaws for up to 5 locations before and after the reference frame are predicted, and then the median is taken out of those yaw predictions.
For stationary instances, we adapt the 3D box fitting algorithm of Zhang et al. [37]. The algorithm first flattens all points into BEV, and then it iterates over all possible angles $\theta\in{0,\frac{\pi}{2}}$. For each angle, two perpendicular axes to each other are computed, and all points are projected into both axes $C_{1}$ and $C_{2}$. Further, the criterion is computed, and $\theta$ with the minimal loss is chosen.
We propose a novel Saturated Closeness Criterion (algorithm 1). Instead of taking a simple sum of distances $d$, as Closeness Criterion in Zhang et al. [37], we feed the $d$ into a sigmoid multiplied by steepness parameter $\alpha$ (eq. 10), inspired by Template Fitting Loss in [31], to saturate the distance of outliers and reduce their influence on the fitting. As a result, the robustness and accuracy of estimation in noisy point clouds are significantly improved over the original algorithm.
$d$ $=\sigma(\alpha\cdot d)$ (10)
Another modification is that Closeness Criterion computes the distance to either minimum or maximum. However, those extremal points tend to be outliers. To mitigate the problem, each point is assigned either to the 10th or 90th percentile of the projected points instead of min/max. Intuitively, this creates an axis at the most probable location, mitigating outliers’ influence.
Algorithm 1 Saturated Closeness Criterion
1:$k$triangleright Steepness parameter
2:function SaturatedCloseness($C_{1},C_{2}$)
3:     $c_{1}^{\text{max}}\leftarrow P_{90}(C_{1})$, $c_{1}^{\text{min}}\leftarrow P_{10}(C_{1})$triangleright P is percentile
4:     $c_{2}^{\text{max}}\leftarrow P_{90}(C_{2})$, $c_{2}^{\text{min}}\leftarrow P_{10}(C_{2})$
5:     $D_{1}\leftarrow\arg\min_{v\in\{c_{1}^{\text{max}}-C_{1},C_{1}-c_{1}^{\text{min}}\}}\|v\|_{2}$
6:     $D_{2}\leftarrow\arg\min_{v\in\{c_{2}^{\text{max}}-C_{2},C_{2}-c_{2}^{\text{min}}\}}\|v\|_{2}$
7:     $D_{1}\leftarrow\sigma(\alpha\cdot D_{1})$triangleright $\sigma$ is the sigmoid function
8:     $D_{2}\leftarrow\sigma(\alpha\cdot D_{2})$
9:     $L\leftarrow 0$
10:     for  $i=1$ to $\text{length}(D_{1})$  do
11:         $d\leftarrow\min(D_{1}(i),D_{2}(i))$
12:         $L\leftarrow L+d$
13:     end for
14:     return $L$
15:end function
As mentioned, the algorithm chooses the yaw with a minimal loss. It creates two hypotheses, as it doesn’t differentiate between the front and back of the car. It is worth noting that the novel Saturated Closeness Criterion does not significantly increase the computational requirements.
To make the estimation more robust, we take both the $A_{j}$ and multiple $P_{i,j}$. Our method chooses frames so the object instance is the closest possible to the ego-vehicle, but at the same time, it is still fully in view to benefit from the fact that the pseudo-lidar is the most accurate for close predictions.
The algorithm outputs a single spatial dimension for a single frame. If any of the values exceed the expected dimensions of an instance, the output is replaced by a prior estimate of the dimensions of a generic instance. Also, it is necessary to take into account that estimating the spatial dimensions of an instance from a single image can be an ill-posed problem; for example, if an instance is seen from the back, it is not possible to estimate length correctly. Thus, such cases are detected, and again, the output is replaced by a prior car dimensions estimate.
Position refinement
The yaw $\theta_{i,j}$ and an approximate 3D position $L_{i,j}$ are already known from the previous steps. In order to get a more fine-grained estimate of position $L_{i,j}$, we apply small perturbations (up to 2 meters) along $x$ and $z$ axes and use Template Fitting Loss (TFL) [31] as the criterion to select the final position. In this stage, we also determine whether the car is facing towards or away from the ego-vehicle by testing two hypotheses for vehicle orientation – $\theta$ and $\theta+\pi$ and again selecting the orientation with the lowest TFL. As the orientation, location and dimensions are known at this step, 3D pseudo labels are created and further used to train the Monocular 3D Detector.
III-B Canonical Object Space (COS)
Figure 5: Inherent ambiguity caused by different camera focal length. The same scene (left) captured by two cameras with different focal length (middle) results in two images (right) where the object has different size. A monocular 3D object detector is however expected to predict the same 3D object location for both images, which is impossible without the detector knowing the focal length as well.
Camera focal length is a crucial parameter for precise metric depth estimation, but monocular 3D detectors, however, have no direct way to compensate for changing focal length because the same object at the same distance from the camera will have different size in the image, depending on the focal length (see fig. 5).
Assuming an object of size $S^{\text{real}}$ with a distance $D$ from the camera in world coordinates, the apparent size ($\approx$ number of pixels) in the image $S^{\text{img}}$ is given as
$\frac{S^{\text{real}}\cdot f_{1}}{D}=S^{\text{img}}_{1}\neq S^{\text{img}}_{2}=\frac{S^{\text{real}}\cdot f_{2}}{D}$ (11)
where $f_{1}$ and $f_{2}$ denote the focal length of camera capturing the same scene.
Taken to the extreme, for the exact same object image (same size $S^{\text{img}}$), the network needs to predict arbitrarily different distances in world coordinates due to the focal length change. This is clearly impossible, and this phenomenon confuses the network both in training as well as at inference time. When combining multiple datasets that were not taken with a single camera or even a single dataset that used different cameras to capture the dataset [38], it is a pre-requisite to be able to accommodate focal length differences.
Figure 6: Canonical Object Space. One scene viewed with two different cameras with different parameters. As the imaging size of the same object is different, 3D labels are scaled accordingly to adapt for focal length change. The predictions and loss computations are all calculated in the Canonical Object Space, and the final predictions are obtained after scaling back based on (known) camera parameters.
Inspired by Metric3D [3] where depth maps are normalized into a uniform space, our method presents Canonical Object Space (COS) where3D bounding boxes are trained to be invariant to changes in the focal length by choosing a single canonical focal length and then transforming the labels (objects) into COS. Instead of transforming whole images or depth maps, as in Metric3D, only the $x$,$y$,$z$ coordinates of all pseudo-ground truth labels are transformed. To transform labels, the scaling parameter $\omega_{i}$ is calculated as
$\omega_{i}=\frac{f^{C}}{f_{i}}$ (12)
where $f_{i}$ stands for the focal length of the frame $i$ and $f^{C}$ is the canonical focal length. Given the scaling parameter $\omega_{i}$, the object instance $(x,y,z)$ is transformed into COS as
$x^{C}=x\cdot\omega_{i}\qquad y^{C}=y\cdot\omega_{i}\qquad z^{C}=y\cdot\omega_{i}$ (13)
The model then operates in the COS and is directly supervised by the transformed labels (objects). During inference, all predictions are transformed from COS by dividing the depth prediction by scaling parameter $\omega_{j}$ of the inference frame $j$
$x=\frac{x^{C}}{\omega_{j}}\qquad y=\frac{y^{C}}{\omega_{j}}\qquad z=\frac{z^{C}}{\omega_{j}}$ (14)
This very simple transformation enables the network to train on multiple datasets where focal length differs and also allows the network to work in unseen camera setup (focal length) because $\omega_{j}$ is calculated at inference time (see table III – 0% fine-tuning row).
Note that in the training process, it is necessary to adapt for data augmentations that affect the perceived focal length of the image, such as image scaling, and for the fact that the detector resizes the image into a constant input size by adjusting the perceived focal length of the image subject to the applied augmentations and the resizing.
IV Experiments
Datasets
We use two public datasets for the evaluation – KITTI [8] and KITTI-360 [39]. KITTI dataset setup has four front-facing cameras, IMU and LiDAR. KITTI-360 setup uses two front-facing cameras and two fish-eye cameras facing sides, IMU and LiDAR. The cameras and their parameters differ in both datasets. For the KITTI dataset, the same training (3712 samples) and validation (3769 samples) splits as [2, 40, 29, 31] are used. For the KITTI-360 dataset, the training (6 sequences, 44178 frames), validation (2 sequences, 1115 frames) and testing (1 sequence, 2459 frames) splits from VSRD [4] are used. It is worth noting that, unlike VSRD, we are able to use all frames of the training set as our method is much faster.
Implementation
We used off-the-shelf MViT2-Huge [41] in the Detectron2 framework [42] trained on MS-COCO [43] as the 2D object detector and Metric3DV2-giant [9] for metric depth estimation. We use our pseudo-labels to train MonoDETR [19] monocular 3D object detector, using AdamW [44] optimiser with learning rate and weight decay equal to 0.0002 and 0.0001, respectively, while keeping other hyper-parameters as in  [19]. The canonical focal length is equal to 500. Aggregation is done over 100 frames, the $p$-score threshold for stationary/moving classification is 0.0001, and the minimum threshold for $d_{real}$ is 5 meters. The steepness parameter $\alpha$ in the Saturated Closeness Criterion is 10.
IV-A Results
Training supervision 3D Object detection accuracy Labelling
Human Labels APBEV/AP3D@0.5 APBEV/AP3D@0.3 speed
Method LiDAR Masks 3D boxes Easy Hard Easy Hard per frame
Fully-supervised methods
MonoFlex [23] yes* yes 50.82/43.11 41.78/34.43 69.70/67.07 59.86/57.26 -
MonoDETR [19] yes* yes 47.21/41.01 36.05/30.38 63.07/60.49 54.04/50.03 -
Methods requiring some human annotations
WeakM3D [1] yes yes 8.10/2.96 2.96/2.01 29.89/21.25 24.01/15.34 80 ms
Autolabels [2] yes yes 20.18/4.69 14.33/2.79 48.16/12.92 37.34/9.94 6 sec
VSRD [4] yes 29.07/21.77 22.83/16.46 58.40/50.86 50.61/43.45 15 min
Methods without human annotations
Ours 38.41/29.98 35.26/27.56 50.84/42.72 49.22/46.59 5 sec
TABLE I: Evaluation results of monocular 3D object (vehicle) detection on the KITTI-360 test set. * denotes that given data modality was not used by the method itself, but it was required in the annotation process to create human 3D box annotations.
Dataset Labels APBEV/AP3D@0.5 APBEV/AP3D@0.3
for training Easy Moderate Hard Easy Moderate Hard
KITTI pseudo 59.76/51.55 44.08/37.09 36.99/33.15 73.38/72.70 57.23/56.30 48.59/47.70
K360 pseudo 57.62/47.39 43.86/38.33 37.05/32.23 75.83/74.40 63.69/61.56 55.75/53.85
KITTI + K360 pseudo 62.28/53.04 46.38/40.79 42.38/35.10 79.70/78.48 65.85/63.97 57.41/55.46
KITTI human 67.44/65.09 53.46/47.39 46.80/44.58 80.30/79.72 67.16/65.87 59.54/58.83
TABLE II: Monocular 3D object detection accuracy on the KITTI validation set, depending on which datasets/labels are used in training.
KITTI-360.
The first evaluation and comparison with prior work is on KITTI-360 [39] as shown in table I. Our method shows outstanding results on both the Bird’s eye view (BEV) and 3D at 0.5 IoU as it outperforms current state-of-the-art weakly-supervised method VSRD [4], despite the fact our method does not use any human labels.
In the 0.3 IoU metric both for BEV and 3D, our method outperforms both WeakM3D [1] and Autolabels [2] by a significant margin and it achieves a competitive AP as VSRD [4]. We believe that this is likely caused by the detection failures, class ambiguities or that KITTI 2D bounding boxes are amodal as using consistent ground-truth 2D labels significantly helps VSRD.
Cross-dataset evaluation
Thanks to the proposed Canonical Object Space, it is possible to train on multiple merged datasets. We, therefore, combine both KITTI and KITTI-360 datasets training subsets, generate our pseudo-labels, train a single model and evaluate it on the KITTI validation subset (see table II). We show that generating pseudo-labels for two different datasets and then performing training on the joint significantly improves average precision both for 0.3 and 0.5 IoU on BEV and 3D. The gain comes from the additional amount of data present in the KITTI-360 dataset.
We also demonstrate the ability of our method to work out-of-the-box in previously unseen camera setup (see table II - row K360), where the model was trained using pseudo-labels on KITTI-360 but evaluated on in this case unseen KITTI dataset. We note that considering IoU 0.3, the accuracy of a model trained on KITTI-360 is actually even slightly higher than the model trained on the original domain (KITTI), which we contribute simply to the fact that KITTI-360 contains more training data.
Semi-supervised
GT APBEV/AP3D@0.7
Method Ratio Easy Moderate Hard
VSRD [4] 0% 0.002/0.001 0.004/0.001 0.005/0.002
Ours 24.46/8.24 20.03/6.84 16.14/6.36
VSRD [4] 25% 31.72/21.76 22.32/15.43 18.86/12.55
Ours 39.99/32.64 30.92/25.06 26.47/20.35
VSRD [4] 50% 43.44/31.05 31.54/21.48 27.17/17.93
Ours 44.32/33.70 32.97/25.65 27.52/23.08
VSRD [4] 75% 42.58/32.95 31.08/24.68 27.19/21.38
Ours 43.32/34.11 32.48/25.68 27.39/23.05
MonoDETR [19] 100% 37.99/29.36 26.76/20.64 23.02/17.30
TABLE III: 3D object detection accuracy on KITTI validation set depending on the fraction of human labels used to fine-tune the model.
In this experiment, MonoDETR is trained jointly on KITTI and KITTI-360, while all labels are generated by our auto-labelling pipeline, as a pre-training step. Next, only a given fraction of human labels is used to fine-tune the model.
Firstly, as seen in table III, thanks to COS our method works out-of-the-box without any fine-tuning (0% human labels), unlike previous methods, showing there is no need to fine-tune using human labels while moving to different domains. Secondly, pre-training on a larger dataset auto-labelled by our method and then fine-tuning on the target domain significantly improves the average precision over simply training on the target domain from scratch. With only 25% of the ground truth labels, our method outperforms traditional fully-supervised detector trained from scratch with 100% of human labels, and the gap grows as more human labels are used. Lastly, our method significantly outperforms current state-of-the-art VSRD [4] by a large margin, showing that pre-training on larger datasets is a promising direction for further research.
Figure 7: Example 3D car detections of our model trained without using human annotations on both KITTI [8] and KITTI-360 [39].
IV-B Ablations
Dataset APBEV/AP3D@0.3
for training COS Easy Moderate Hard
KITTI 80.30/79.72 67.16/65.87 59.54/58.83
KITTI+K360 71.71/69.62 59.44/57.89 53.76/52.25
KITTI+K360 82.07/81.25 67.14/66.16 62.47/61.02
TABLE IV: Canonical Object Space (COS) ablation on KITTI validation set under full supervision from human labels.
Canonical Object Space
As seen in table IV, disabling COS and training on both datasets jointly decreases the average precision of both BEV and 3D. It may be surprising that when training on both KITTI+K360, the model is not broken completely – we speculate that the network is able to learn to classify from which of the two datasets the image comes from and adjusts its outputs accordingly; such an approach obviously does not generalize well. Note that only in this ablation, we use human labels to eliminate any effects of pseudo-labels.
Number of frames used in aggregation
No. of APBEV/AP3D@0.5
frames Easy Moderate Hard
$\pm$ 1 53.26/49.35 41.34/35.86 34.77/29.14
$\pm$ 10 59.76/61.55 44.08/37.09 36.99/33.15
$\pm$ 30 59.22/51.18 42.65/36.64 35.91/29.77
TABLE V: Number of frames used in the aggregation process ablation on KITTI validation set.
table V shows the ablation study on the number of frames used in the aggregation process. Perhaps surprisingly, aggregating over longer sequences does not yield better performance. The aggregation can recover from the error caused by Metric3D [3, 9] as in fig. 4. We speculate that the distant car poses a significant challenge for the detector itself, therefore creating precise 3D labels for such objects does not translate to better 3D detection accuracy.
APBEV/AP3D@0.5
$\mathbf{\alpha}$ Easy Moderate Hard
1 53.02/48.09 41.29/34.92 34.48/28.66
5 55.14/50.83 42.46/36.59 35.92/29.85
10 56.56/48.90 41.32/35.00 34.53/28.76
15 55.78/51.86 42.48/36.63 35.72/29.81
TABLE VI: Steepness parameter $\alpha$ of the Saturated Closeness Criterion ablation on KITTI validation set.
Steepness parameter in Saturated Closeness Criterion
In this ablation, we explore possible values for the steepness parameter $\alpha$ in the Saturated Closeness Criterion, which controls what pseudo-LIDAR points are discarded as outliers. As shown in table VI, $\alpha$ equal to 10 achieves the best performance.
V Conclusion
A novel method for training monocular 3D detectors without domain-specific human annotations was presented. The method exploits temporal consistency in video sequences to automatically create 3D labels of objects (cars) and, as such, does not require human labels or additional sensors such as LiDAR. The method is able to compensate for camera focal length differences to significantly improve its scalability across multiple datasets, and furthermore, it works out of the box in previously unseen camera setups.
The method is evaluated on two large-scale public datasets, where despite using no human labels, it outperforms prior work by a significant margin both for BEV and 3D when using the stricter 0.5 IoU evaluation. Last but not least, the method shows its versatility by also being a powerful pre-training method for fully-supervised monocular 3D detection, where the final accuracy increases by approximately 6 percent points over traditional object detector.
The main limitation is detecting objects far away, which is not primarily due to the auto-labelling procedure but due to the inherent limitation of the trained monocular detector to infer the distance of objects which are only a couple of pixels high.

### v2

@@
 Abstract
-Detecting the three-dimensional position and orientation of objects using a single RGB camera is a foundational task in computer vision with many important applications.
-Traditionally, 3D object detection methods are trained in a fully-supervised setup, requiring vast amounts of human annotations, which are laborious, costly, and do not scale well with the ever-increasing amounts of data being captured.
-In this paper, we present the first method to train 3D object detectors for monocular RGB cameras without domain-specific human annotations, thus making orders of magnitude more data available for training.
-Thanks to newly proposed Canonical Object Space, the method can not only exploit data across a variety of datasets and camera setups to train a single 3D detector, but unlike previous work it also works out of the box in previously unseen camera setups.
-All this is crucial for practical applications, where the data and cameras are extremely heterogeneous.
-The method is evaluated on two standard autonomous driving datasets, where it outperforms previous works, which, unlike our method, still rely on 2D human annotations.
-The source code and the model is going to be published soon.
+Inferring object 3D position and orientation from a single RGB camera is a foundational task in computer vision with many important applications.
+Traditionally, 3D object detection methods are trained in a fully-supervised setup, requiring LiDAR and vast amounts of human annotations, which are laborious, costly, and do not scale well with the ever-increasing amounts of data being captured.
+We present a novel method to train a 3D object detector from a single RGB camera without domain-specific human annotations, making orders of magnitude more data available for training.
+The method uses newly proposed Local Object Motion Model to disentangle object movement source between subsequent frames, is approximately 700 times faster than previous work and compensates camera focal length differences to aggregate multiple datasets.
+The method is evaluated on three public datasets, where despite using no human labels, it outperforms prior work by a significant margin.
+It also shows its versatility as a pre-training tool for fully-supervised training and shows that combining pseudo-labels from multiple datasets can achieve comparable accuracy to using human labels from a single dataset.
+The source code and model will be published soon.
 (a) Traditional pipelines require human labelling and produce new model for every different camera setup
@@
 Monocular 3D object detection is a key component of many important systems, ranging from robotics to autonomous cars.
-3D object detection methods are typically trained in a fully-supervised setup, requiring vast amounts of human annotations, which are laborious, costly, and do not scale well with ever-increasing amounts of data that ideally should be used for training.
-To make matters even worse, especially in autonomous driving, new cameras and new camera setups emerge constantly, which typically implies the need to collect, label and train on a new dataset as camera parameters such as focal length change.
-In this paper, we address both of these challenges – costly human labels and changing camera setup – by introducing a new method to train a 3D rigid object detector in a Canonical Object Space without domain-specific human labels.
-Our method allows us to exploit large amounts of training data readily available because all we need as input for training are unlabelled driving video sequences.
-Furthermore, our method does not require data from additional sensors such as LiDAR.
-It makes our method extremely applicable in practice because as such our method can directly exploit data collected by a majority of currently manufactured vehicles, which typically only have a single camera.
-This is in contrast with some methods which require LiDAR (and human annotations) [1, 2], because for them specialised vehicles are needed to collect the data, and this severely limits the variety of geographies, environments, and traffic situations being captured.
-Drawing inspiration from monocular depth literature [3], the proposed Canonical Object Space (COS) allows us to aggregate and learn from data coming from different cameras, which is a crucial pre-condition to being able to effectively harvest otherwise very heterogeneous video sequences.
-Moreover, unlike previous methods such as VSRD [4], it also allows for the deployment of our 3D object detector in new unseen camera setup without needing to train or fine-tune for the new setup.
-This is again extremely useful in practice, because new camera sensors and setups are developed all the time, but it is unfeasible to collect data, label them by human annotators, and then train a new model for each new camera type.
+3D object detection methods are typically trained in a fully-supervised setup, requiring specialized vehicles to simultaneously capture RGB image and LiDAR modalities, followed by human annotators manually creating 3D labels.
+This is extremely laborious, costly, and limits the diversity of training data being captured.
+To make matters even worse, especially in autonomous driving, new camera setups emerge constantly, which typically means this laborious and costly process needs to be repeated again.
+In this paper, we address both of these challenges – costly human labelling and reliance on dedicated vehicles to capture data – by introducing a new method to train a monocular 3D rigid object detector without domain-specific human labels and without LiDAR.
+Our method allows us to exploit large amounts of training data readily available because it does not rely on human annotations, which traditionally is the main limiting factor in exploiting captured data, and it does not need additional sensors such as LiDAR, which implies our method can directly exploit data collected by a majority of currently manufactured vehicles, which typically only have a single camera.
+This is in contrast with some methods which require LiDAR (and human annotations) [1, 2] as specialized vehicles are needed to collect the data.
+This then severely limits the diversity of geographies, environments, and traffic situations captured.
+In our method, we automatically create 3D object labels (7-DOF poses) by using outputs of two generic off-the-shelf models – a 2D object detector and a monocular depth estimator – and by combining them into a single pipeline that processes video sequences captured from a moving vehicle.
+The proposed pipeline uses temporal consistency of objects in real world to precisely infer object 3D object orientation, position and dimensions in a series of consecutive steps, using newly proposed Local Object Motion Model (LOMM).
+LOMM disentangles the position change between subsequent frames into ego-motion of the vehicle capturing the scene and motion of the labeled object itself, allowing our method, unlike previous work [3, 4], to aggregate information for both stationary and moving objects throughout the whole video sequence, which then results in more precise 3D orientation and position 3D label in every frame.
+To address the heterogeneity of captured data in practice, we are the first weakly-supervised method to address this issue by adopting Canonical Object Space (COS) [5, 6] to effectively harvest and auto-label video sequences captured by different camera setups.
 To summarize, we make the following contributions:
   1. 1.
-    We propose a novel auto-labelling pipeline which exploits temporal consistency in video sequences to automatically generate 3D position, size, and orientation labels of rigid objects (cars) from a monocular camera.
+    To the best of our knowledge, our method is the first method to train a monocular 3D object detector for autonomous driving without 2D or 3D human annotations; we demonstrate to the community that such an approach is feasible and is worth pursuing further.
   2. 2.
-    To the best of our knowledge, our method is the first method to train a monocular 3D object detector for autonomous driving which does not require neither 2D nor 3D human annotations; we demonstrate to the community that such an approach is feasible and is worth pursuing further.
+    We propose a novel auto-labelling pipeline which exploits temporal consistency in video sequences to generate 3D position, size, and orientation labels of rigid objects (cars) from a monocular camera.
+The pipeline uses novel Local Object Motion Model (LOMM) to disentangle ego-motion of the vehicle capturing the scene and motion of the labelled object itself to aggregate information both for stationary and moving objects.
   3. 3.
-    We propose novel Canonical Object Space (COS), which allows us to a) aggregate data from different camera setups in training, and b) at inference time, use an already trained model in the previously unseen camera setup, i.e.
-without any training or fine-tuning for the new setup.
-  4. 4.
-    Our auto-labelling method is significantly faster than previous work, allowing us to process significantly more training data than the current state-of-the-art method.
+    Our auto-labelling method is almost 700 times faster than previous work and is able to aggregate data captured by different camera setups, allowing us to process significantly more data than the current state-of-the-art methods [4], where their auto-labelling speed in practice prohibits applications to large-scale datasets.
 II Related work
-Monocular Depth Estimation
+Monocular Depth Estimation.
 Monocular depth estimation has shown significant improvements in recent years.
@@
 Depth Anything [9] focuses on the affine-invariant depth estimation using self-supervised pre-training.
-For metric depth estimation, which is relevant for object detection, it fine-tunes on the target dataset such as KITTI [8].
+For metric depth estimation, which is relevant for 3D object detection, it fine-tunes on the target dataset such as KITTI [10].
 Metric3D [11, 5] focuses on metric depth estimation as it uses a Canonical Camera Transformation Module to further improve zero-shot performance, together with DINOv2 [12] or ConvNext [13] as the backbone and DPT [14] as the predictor.
-CamConv [13] models cameras inside the network, but as shown in [3], this leads to poor results.
-Fully-supervised Monocular 3D Object Detection
+We opt to use Metric3D [11, 5] for its superior zero-shot performance.
+Fully-supervised Monocular 3D Object Detection.
 Early methods [15, 16, 17] were based on lifting the image into pseudo-LiDAR and then detecting objects in the generated point cloud.
@@
 MonoCD [25] enhances the ensemble with a complementary depth predictor.
-However, all the aforementioned methods require 3D human annotations during their training.
-Weakly-supervised 3D Object Detection
-Most previous works on weakly-supervised 3D detection have focused on training 3D detectors for LiDAR.
-VS3D [25] pioneered in this area by using point density in LiDAR scans to identify objects, and detections are further refined by a student-teacher network, where the teacher is trained on images only.
+However, all the aforementioned methods were previously trained using 3D human labels.
+We utilize MonoDETR [20] for its competitive performance and high-quality code implementation.
+Weakly-supervised LiDAR 3D Object Detection.
+The majority of previous work on weakly-supervised 3D detection have focused on training 3D detectors for LiDAR.
+VS3D [26] pioneered in this area by using point density in LiDAR scans to identify objects.
 Zakharov et al.
-[2] employs a 2D off-the-shelf detector accompanied by a novel Signed Distance Fields (SDF)-renderer within the DeepSDF framework [26] and Normalized Object Coordinate Space (NOCS).
-The method is trained on a synthetic dataset and further fine-tuned on real datasets.
-WS3D [27, 28] employs Bird’s Eye View (BEV) human-centered clicks as weak supervision to train a network for detecting vehicles, arguing that these are much easier to obtain than 3D bounding boxes.
-Furthermore, they use a small subset of human ground-truth annotations to train a second network to predict the spatial dimensions of the detected vehicles.
+[2] employs a 2D off-the-shelf detector accompanied by a novel Signed Distance Fields (SDF)-renderer and Normalized Object Coordinate Space (NOCS).
+WS3D [27, 28] employs Bird’s Eye View (BEV) human-centred clicks as weak supervision to train a network for detecting vehicles.
 McCraith et al.
-[29] employ direct fitting of generic templates into LiDAR scans segmented by Mask-RCNN [30] and share the information between frames by requiring consistency between frames while simultaneously discounting outliers.
-TCC-det [31] uses direct fitting of generic templates on aggregated points corresponding to each instance.
-During the training of the 3D detector, TCC-det appends two additional losses for finer alignment.
-VG-W3D [32] encodes the image and LiDAR scan separately and then enforces alignment between 2D and 3D on multiple feature levels to predict precise 3D bounding boxes.
-Figure 2: Our auto-labelling pipeline creates 3D pseudo-groundtruth using only images, camera parameters and IMU data as inputs.
-In contrast to the aforementioned methods, VSRD [4] focuses on 3D detection in camera, and does not require LiDAR scans.
-However, it uses 2D human ground-truth masks for each instance, representing them as a surface in Signed Distance Fields (SDF).
-Furthermore, the SDF are rendered into masks through their proposed volumetric-aware silhouette rendering, enabling end-to-end training.
+[29] employ direct fitting of generic templates into LiDAR scans while requiring consistency between frames for stationary vehicles and discounting outliers.
+TCC-det [30] uses direct fitting of generic templates on aggregated information corresponding to each instance and trains with two additional losses for finer alignment.
+VG-W3D [31] encodes the image and LiDAR scan separately and then enforces alignment between 2D and 3D on multiple feature levels to predict precise 3D boxes.
+Figure 2: Our auto-labelling pipeline creates 3D labels using only images and ego-motion data as inputs (and known camera parameters).
+Unsupervised LiDAR 3D Object Detection
+Several approaches have been proposed for unsupervised object detection and tracking in LiDAR scans.
+Cheng et al.
+[32] focus on open-set 3D object detection using the MLUC network, which identifies objects while keeping some unclassified.
+ClusterNet [33] performs unsupervised object discovery using both 3D LiDAR and 2D RGB images.
+Najibi et al.
+[34] involve scene-flow estimation and meta-label creation for detection and trajectory prediction.
+MODEST [35] detects objects via clustering and refines them in a self-training cycle.
+DRIFT [36] uses a Persistency prior score for noisy label creation and reinforced learning for fine-tuning.
+Recent approaches include density-based clustering and self-supervised learning for object detection [37], clustering based on the ”what moves together” paradigm [38], and self-supervised LiDAR flow estimation for motion cue generation [39].
+While [32, 37] do not differentiate between moving and stationary vehicles, [33, 34, 35, 36, 38, 39] only detect moving vehicles and assume the detector will generalize to detect all movable vehicles (including stationary).
+Weakly-supervised Monocular 3D Object Detection
+WeakMono3D [3] is the first method not using LiDAR scans, as only sequences of images with 2D ground-truth labels are required as input.
+VSRD [4] represents each instance as a surface in Signed Distance Fields (SDF).
+The SDFs are rendered into masks through their proposed volumetric-aware silhouette rendering, enabling end-to-end training.
 Each instance is optimized over multiple frames.
-While VSRD shows promising results in both weakly-supervised and semi-supervised training, it requires 2D instance masks, and rendering of each frame takes approximately 15 minutes.
-In contrast, our method does not require any human annotations at all and is approximately 100 times faster.
+Both WeakMono3D and VSRD struggle with moving vehicles, which are either discarded or given low confidence; on the other hand, our work can leverage information from both moving and stationary vehicles.
+While WeakMono3D and VSRD show promising results in both weakly-supervised and semi-supervised training, they require 2D instance masks, and in VSRD rendering of each frame takes approximately 15 minutes.
+In contrast, our method does not require any human annotations at all and is approximately 700 times faster.
 III Method
-Traditionally, deep networks that predict 3D bounding boxes of vehicles (and other agents) from a single image are trained using 3D annotations, which are manually created by human annotators.
-They typically rely on both camera image and LiDAR data for correct distance estimate in creating their labels.
-However, this process is very laborious and costly, limiting the amount of training data available to train such methods.
-Moreover, it requires capturing the scene using both a camera and a LiDAR sensor, which is a significant limitation since the majority of production vehicles these days do not have LiDAR on board – as a result, data collection is only limited to vehicles dedicated to such a purpose, and this again limits the variety and diversity of data being used to train these methods.
-Our method overcomes these limitations by introducing a novel approach to train a 3D object detector using only images (and vehicle motion data) as inputs, without requiring human labels or LiDAR in the training process.
-III-A Auto-labelling pipeline
 The process of automatic label creation assumes that images are captured in a sequence, which is a natural way to capture data in driving scenarios.
-It also assumes that the camera setup is known (both intrinsics and extrinsics) for each sequence.
-However, the camera setup might be and actually often is different (see section III-B).
+It also assumes that the camera setup is known (both intrinsics and extrinsics) for each sequence, although each sequence might have different camera setup.
 Lastly, it assumes approximate data about ego-vehicle’s motion is available (e.g.
-from GPS or an Inertial Measurement Unit – IMU), which again is commonplace.
-Because labelling a 7-DOF pose of objects jointly is a very challenging problem, we instead infer individual pose parameters separately in a sequence of steps.
-Detecting objects in pseudo-LiDAR
+from GPS or IMU), which again is commonplace.
+Because labelling a 7-DOF pose of objects jointly (position, size, yaw) is a very challenging problem, we instead infer individual pose parameters separately in a sequence of steps (see fig.
+2).
+III-A Pseudo-LiDAR aggregation
 The auto-labelling process begins by inferring a metric depth map $D\in\mathbb{R}^{h\times w}$ for each camera image $I\in\mathbb{R}^{3\times h\times w}$, where $h$ and $w$ denotes image height and width, using an off-the-shelf monocular depth estimator.
-We opted to use Metric3D [3], as it is zero-shot, it does not limit our method in terms of data distribution and has shown great results for generalization.
-The choice of a particular monocular metric depth estimator is, however, not crucial.
-For each pixel and its corresponding depth, a 3D point cloud is generated as
-$X_{u,v}$ $=\frac{D_{u,v}\cdot(u-c_{x})}{f_{x}}$
-$Y_{u,v}$ $=\frac{D_{u,v}\cdot(v-c_{y})}{f_{y}}$
+We opted to use Metric3D [5], as it is zero-shot, it does not limit our method in terms of data distribution and it has shown great generalization results.
+Using the inferred depth map $D$, a 3D point cloud is generated as
+$X_{u,v}$ $=\frac{D_{u,v}\cdot(u-c_{x})}{f_{x}}\qquad Y_{u,v}=\frac{D_{u,v}\cdot(v-c_{y})}{f_{y}}$
 $Z_{u,v}$ $=D_{u,v}$ (1)
 where $u\in[0,h]$ and $v\in[0,w]$, $c$ is the principal point and $f$ is the focal length.
-We denote the pseudo lidar point cloud as $P_{i}\in\mathbb{R}^{3\times hw}$, where $i$ stands for the frame index.
-Figure 3: Combining instance segmentation in image (top) with monocular metric depth estimation (middle) enables us to produce pseudo-LiDAR point cloud for each object.
-Next, inspired by [29], we employ an off-the-shelf 2D object image detector to detect the objects we are interested in (vehicles) and their instance segmentation masks in each camera image.
+We denote the pseudo-LiDAR point cloud as $P_{i}\in\mathbb{R}^{3\times hw}$, where $i$ stands for the frame index.
+Figure 3: Combining instance segmentation in image (top) with monocular metric depth estimation (middle) enables us to produce pseudo-LiDAR point cloud for each object (bottom).
+Next, inspired by [29], we employ an off-the-shelf 2D object image detector [40] to detect the objects we are interested in (vehicles) and their instance segmentation masks in each camera image.
 Given an instance segmentation mask $M_{i,j}\in\mathbb{R}^{h\times w}$ of an object $j$ in a frame $i$, we find object point cloud $P_{i,j}$ in the point cloud as
 $P_{i,j}=\big{\{}p\in P_{i}\,|\,\mathcal{T}(p)\in M_{i,j}\big{\}}$ (2)
-where $\mathcal{T}(p)$ denotes the projection of the 3D point $p$ into image coordinates (using camera calibration, which is known).
-The $P_{i,j}$ tends to contain outliers, as either the masks or the metric depth estimation is not perfect.
-To remove the outliers, an ensemble of the following methods is employed: HBOS [33], Z-scores, statistical outlier rejection in Open3D [34], HDBSCAN [35] and DBSCAN [36].
-If at least two of these methods classify a point as an outlier, it is removed.
-To compensate for the effect of the ego-vehicle movement, we transform all $P_{i,j}$ into the reference frame coordinate space, given the poses from IMU data.
-Temporal consistency
-In order to exploit temporal consistency across multiple frames of the sequence, we need to first establish correspondences of individual object instances between frames.
-Because we have inferred approximate 3D point cloud of each object instance $P_{i,j}$ in each frame, we track each object in 3D world coordinate system.
-For each object $j$ in frame $i$, we approximate its location $L_{i,j}$ as the median of its point cloud $P_{i,j}$.
+where $\mathcal{T}(p)$ denotes the projection of the 3D point $p$ into image coordinates, as shown in fig. 3.
+Tracking.
+In order to exploit temporal consistency of given object across the sequence frames, we need first to establish correspondences of the object instance between frames.
+Because we have inferred the approximate 3D point cloud of each object instance $P_{i,j}$ in each frame, we track each object in the 3D world coordinate system.
+We approximate spatial location $L_{i,j}\in\mathbb{R}^{3}$ of each object $j$ in frame $i$ as the median of its point cloud $P_{i,j}$.
 The tracking is initialized in $n$ frames before the reference (current) frame, and for each frame, it matches the instances based on locations in the current frame $L_{i,j}$ and the predicted future locations from the last frame into the current frame $\hat{L}_{i,j}$.
@@
 After this step, each object instance is represented by extracted points $P_{i,j}$ and locations $L_{i,j}$ in each frame the instance is present.
-The tracking sequentially processes all the frames up to $n$ after the reference frame.
-Movement Classification
-The next step is to classify instances as either stationary or moving, as different principles of temporal consistency exploitation are applied to those classes.
-Note that all object instances actually appear moving – they change location in relative terms, as the ego-vehicle itself is driving while data is captured.
+The tracking sequentially processes all the frames up to $m$ after the reference frame.
+Finally, we transform all $P_{i,j}$ into the reference frame coordinate space (current frame for which labels are generated) using vehicle ego-motion data.
+III-B Local Object Motion Model
+In our method, we classify object instances as either stationary or moving, as different principles of temporal consistency exploitation are applied to those classes.
+Note that all object instances actually appear to change position between frames – they change their location in relative terms, as the ego-vehicle itself is driving while data is captured.
 In order to classify a vehicle as stationary or moving, we propose the following novel classification procedure.
-For each instance, all locations $L_{m...n,j}$ are taken and differences between all adjacent locations are computed, where $m$ is the first frame, where the instance is present and $n$ is the last.
+For each instance, all locations $L_{m...n,j}$ are taken, and differences $\Delta_{i,j}$ between all adjacent locations are computed, where $k$ is the first frame, where the instance is present and $l$ is the last.
 From those differences, the standard deviation $\sigma$ is calculated as
-$\sigma_{j}$ $=\frac{1}{\sqrt{2}}\sqrt{\frac{1}{n-m}}\sum_{i=m+1}^{n}(\mu_{j}-D_{i,j})^{2}$ (4)
-$\mu_{j}$ $=\frac{1}{n-m}\sum_{i=m+1}^{n}D_{i,j}\qquad D_{i,j}=L_{i,j}-L_{i-1,j}$ (5)
-Furthermore, the actual distance $d$ and expected distance $\hat{d}$ caused by the $\sigma$ is calculated as
-$d_{j}=\left\lVert L_{m,j}-L_{n,j}\right\rVert^{2}_{2}\qquad\hat{d}_{j}=(n-m)\cdot\left\lVert\sigma_{j}\right\rVert^{2}_{2}$ (6)
-The $z$- and $p$-scores are calculated to know if the difference between the $d_{\text{real}}$ and $d_{\text{exp}}$ is statistically significant or not
-$z_{j}=\frac{d_{j}}{\hat{d}_{j}}\qquad p_{j}=1-\Phi(z_{j})$ (7)
-To classify the instance as moving, the $p$-score is required to be lower than the threshold $T_{m}$ and simultaneously, the $d$ distance must be higher than a given threshold.
-As all $P_{i,j}$ are in the reference frame, we can aggregate points for stationary instances by simply concatenating the points
-$A_{j}=\bigcup_{i=m}^{n}P_{i,j}$ (8)
-Aggregating points for moving instances poses a significant challenge, as the $P_{i,j}$ differs significantly between frames.
-For that reason, for moving instances, only the trajectory and $P_{i,j}$ are employed.
+$\sigma_{j}$ $=\frac{1}{\sqrt{2}}\sqrt{\frac{1}{l-k}}\sum_{i=k+1}^{l}(\mu_{j}-\Delta_{i,j})^{2}$ (4)
+$\mu_{j}$ $=\frac{1}{l-k}\sum_{i=k+1}^{l}\Delta_{i,j}\qquad\Delta_{i,j}=L_{i,j}-L_{i-1,j}$ (5)
+Furthermore, to decide if the difference between the $\left\lVert\mu_{j}\right\rVert_{2}$ and $\left\lVert\sigma_{j}\right\rVert_{2}$ is statistically significant or not, the ratio $z$ is calculated as
+$z_{j}=\frac{\left\lVert\mu_{j}\right\rVert_{2}}{\left\lVert\sigma_{j}\right\rVert_{2}}$ (6)
+To classify the instance as moving, the ratio $z$ is required to be higher than the threshold $T_{z}$ and simultaneously, the net distance must be higher than a given threshold $T_{m}$.
+This follows the idea, that if the car is stationary, the $\left\lVert\mu_{j}\right\rVert_{2}$ will be significantly lower than $\left\lVert\sigma_{j}\right\rVert_{2}$ as the movement of the vehicle is introduced by jitter from pseudo-LiDAR.
+On the other hand, if the vehicle is moving the $\left\lVert\mu_{j}\right\rVert_{2}$ will be significantly higher than $\left\lVert\sigma_{j}\right\rVert_{2}$.
+As a result, for moving instances, we estimate the trajectory in adjacent frames.
+Given the known trajectory and the fact that cars move by physics constraints, estimating the yaw (the orientation) of the car is trivial.
+This simple fact simplifies the autolabelling significantly.
+On the other hand, for the stationary instances, we can simply aggregate points $A_{j}$ by concatenating all $P_{i,j}$, as all $P_{i,j}$ are in the reference frame.
+This creates a much denser and more informative representation of an instance, and it can also recover from some mistakes, as shown in fig.
+4.
 Refer to caption Refer to caption
@@
 c) d)
-Figure 4: Temporal aggregation examples.
-The point cloud of a significantly occluded vehicle (a) is recovered in subsequent frames (b) where the vehicle is no longer occluded.
+Figure 4: The point cloud of a significantly occluded vehicle (a) is recovered in subsequent frames (b).
 Distant car with poor depth prediction (c) is refined as the ego-vehicle drives closer (d).
-White points is pseudo-LiDAR of each frame, colourful points are the aggregated points.
-Human ground truth overlaid as green boxes for reference purposes.
-Orientation and size estimation
-Finding oriented 3D bounding boxes is a 7-degrees of freedom (DOF) problem. Instead of solving it directly, our method divides it into three disjunctive problems: orientation, location and dimensions.
+Pseudo-LiDAR of each frame in white, aggregated object points in purple, human label in green for reference.
+III-C Step-wise autolabelling
+Finding oriented 3D bounding boxes is a 7-degrees of freedom (DOF) problem.
+Instead of solving it directly, our method divides it into three disjunctive problems: orientation, location and size estimation.
+Orientation and size estimation.
 Orientation estimation is different for moving and stationary instances. For moving instances, the yaw of the object is estimated by directly computing the angle between adjacent locations as
 $\theta_{i,j}=\operatorname{atan2}\left(\frac{L_{i,j}(z)-L_{i-1,j}(z)}{L_{i,j}(x)-L_{i-1,j}(x)}\right)$ (7)
-It is worth noting again that all locations are transformed into the reference coordinate space.
 To make it more robust, yaws for up to 5 locations before and after the reference frame are predicted, and then the median is taken out of those yaw predictions.
-For stationary instances, we adapt the 3D box fitting algorithm of Zhang et al.
+For stationary instances, we improve the 3D box fitting algorithm of Zhang et al.
 [41].
-The algorithm first flattens all points into BEV, and then it iterates over all possible angles $\theta\in{0,\frac{\pi}{2}}$.
-For each angle, two perpendicular axes to each other are computed, and all points are projected into both axes $C_{1}$ and $C_{2}$.
-Further, the criterion is computed, and $\theta$ with the minimal loss is chosen.
-We propose a novel Saturated Closeness Criterion (algorithm 1).
-Instead of taking a simple sum of distances $d$, as Closeness Criterion in Zhang et al.
-[37], we feed the $d$ into a sigmoid multiplied by steepness parameter $\alpha$ (eq.
-10), inspired by Template Fitting Loss in [31], to saturate the distance of outliers and reduce their influence on the fitting.
+The algorithm first flattens all points into Bird’s eye view (BEV), and then it iterates over all possible angles $\theta\in{0,\frac{\pi}{2}}$.
+For each angle, two perpendicular axes to each other are computed, and all points are projected into both axes $x\in\mathbb{R}$ and $y\in\mathbb{R}$.
+Algorithm 1 Saturated Closeness Criterion
+1:$\alpha$triangleright Steepness parameter
+2:function SaturatedCloseness($x,y$)
+3:    $x^{\text{max}}\leftarrow P_{90}(x)$, $x^{\text{min}}\leftarrow P_{10}(x)$triangleright P is percentile
+4:    $y^{\text{max}}\leftarrow P_{90}(y)$, $y^{\text{min}}\leftarrow P_{10}(y)$
+5:    $x\leftarrow\arg\min_{v\in\{x^{\text{max}}-x,x-x^{\text{min}}\}}\|v\|_{2}$
+6:    $y\leftarrow\arg\min_{v\in\{y^{\text{max}}-y,y-y^{\text{min}}\}}\|v\|_{2}$
+7:    $x\leftarrow\sigma(\alpha\cdot x)$triangleright $\sigma$ is the sigmoid function
+8:    $y\leftarrow\sigma(\alpha\cdot y)$
+9:    $\mathcal{L}\leftarrow 0$
+10:    for  $i=1$ to $\text{length}(x)$  do
+11:        $e\leftarrow\min(x(i),y(i))$
+12:        $\mathcal{L}\leftarrow\mathcal{L}+e$
+13:    end for
+14:    return $\mathcal{L}$
+15:end function
+For each $\theta$, we calculate a newly proposed Saturated Closeness Criterion (algorithm 1).
+Instead of taking a simple sum of distances $x$ and $y$, as Closeness Criterion in Zhang et al.
+[41], we feed the $x$ and $y$ into a sigmoid multiplied by steepness parameter $\alpha$ (Line 6 and 7 of algorithm 1), inspired by Template Fitting Loss in [30], to saturate the distance of outliers and reduce their influence on the fitting.
 As a result, the robustness and accuracy of estimation in noisy point clouds are significantly improved over the original algorithm.
-$d$ $=\sigma(\alpha\cdot d)$ (10)
 Another modification is that Closeness Criterion computes the distance to either minimum or maximum.
@@
 To mitigate the problem, each point is assigned either to the 10th or 90th percentile of the projected points instead of min/max.
+The assignment is based on the minimal distance to one of the extremal points.
 Intuitively, this creates an axis at the most probable location, mitigating outliers’ influence.
-Algorithm 1 Saturated Closeness Criterion
-1:$k$triangleright Steepness parameter
-2:function SaturatedCloseness($C_{1},C_{2}$)
-3:     $c_{1}^{\text{max}}\leftarrow P_{90}(C_{1})$, $c_{1}^{\text{min}}\leftarrow P_{10}(C_{1})$triangleright P is percentile
-4:     $c_{2}^{\text{max}}\leftarrow P_{90}(C_{2})$, $c_{2}^{\text{min}}\leftarrow P_{10}(C_{2})$
-5:     $D_{1}\leftarrow\arg\min_{v\in\{c_{1}^{\text{max}}-C_{1},C_{1}-c_{1}^{\text{min}}\}}\|v\|_{2}$
-6:     $D_{2}\leftarrow\arg\min_{v\in\{c_{2}^{\text{max}}-C_{2},C_{2}-c_{2}^{\text{min}}\}}\|v\|_{2}$
-7:     $D_{1}\leftarrow\sigma(\alpha\cdot D_{1})$triangleright $\sigma$ is the sigmoid function
-8:     $D_{2}\leftarrow\sigma(\alpha\cdot D_{2})$
-9:     $L\leftarrow 0$
-10:     for  $i=1$ to $\text{length}(D_{1})$  do
-11:         $d\leftarrow\min(D_{1}(i),D_{2}(i))$
-12:         $L\leftarrow L+d$
-13:     end for
-14:     return $L$
-15:end function
-As mentioned, the algorithm chooses the yaw with a minimal loss.
-It creates two hypotheses, as it doesn’t differentiate between the front and back of the car.
-It is worth noting that the novel Saturated Closeness Criterion does not significantly increase the computational requirements.
-To make the estimation more robust, we take both the $A_{j}$ and multiple $P_{i,j}$.
-Our method chooses frames so the object instance is the closest possible to the ego-vehicle, but at the same time, it is still fully in view to benefit from the fact that the pseudo-lidar is the most accurate for close predictions.
-The algorithm outputs a single spatial dimension for a single frame.
-If any of the values exceed the expected dimensions of an instance, the output is replaced by a prior estimate of the dimensions of a generic instance.
+The algorithm chooses the yaw with a minimal criterion value and creates two hypotheses, as it doesn’t differentiate between the front and back of the car.
+The algorithm also outputs a spatial dimensions estimate for a single frame.
+If any of the values exceed the typical dimensions of an instance, the output is replaced by a prior estimate of the dimensions of a generic instance.
 Also, it is necessary to take into account that estimating the spatial dimensions of an instance from a single image can be an ill-posed problem; for example, if an instance is seen from the back, it is not possible to estimate length correctly.
-Thus, such cases are detected, and again, the output is replaced by a prior car dimensions estimate.
-Position refinement
-The yaw $\theta_{i,j}$ and an approximate 3D position $L_{i,j}$ are already known from the previous steps.
-In order to get a more fine-grained estimate of position $L_{i,j}$, we apply small perturbations (up to 2 meters) along $x$ and $z$ axes and use Template Fitting Loss (TFL) [31] as the criterion to select the final position.
+Thus, such cases are detected by computing the difference between car orientation and the viewing angle of the car and if the difference is ${0,\frac{\pi}{2},\pi,\frac{3\pi}{2}}$, then the output is replaced by a prior car dimensions estimate.
+Position refinement.
+The yaw $\theta$ and a prior 3D position are already known from the previous steps.
+In order to get a more fine-grained estimate of position, we apply small perturbations (up to 2 meters) along $x$ and $z$ axes and use Template Fitting Loss (TFL) [30] as the criterion to select the final position.
+To compute TFL, we use a generic template mesh of a vehicle sampled into a point cloud.
+Given a generic vehicle point cloud, we move the point cloud according to perturbations and compute TFL at those points.
+The lowest TFL value corresponds to the best fit between those point clouds.
+Unlike Chamfer Distance, it works well with outliers.
 In this stage, we also determine whether the car is facing towards or away from the ego-vehicle by testing two hypotheses for vehicle orientation – $\theta$ and $\theta+\pi$ and again selecting the orientation with the lowest TFL.
-As the orientation, location and dimensions are known at this step, 3D pseudo labels are created and further used to train the Monocular 3D Detector.
-III-B Canonical Object Space (COS)
-Figure 5: Inherent ambiguity caused by different camera focal length.
-The same scene (left) captured by two cameras with different focal length (middle) results in two images (right) where the object has different size.
-A monocular 3D object detector is however expected to predict the same 3D object location for both images, which is impossible without the detector knowing the focal length as well.
-Camera focal length is a crucial parameter for precise metric depth estimation, but monocular 3D detectors, however, have no direct way to compensate for changing focal length because the same object at the same distance from the camera will have different size in the image, depending on the focal length (see fig.
-5).
-Assuming an object of size $S^{\text{real}}$ with a distance $D$ from the camera in world coordinates, the apparent size ($\approx$ number of pixels) in the image $S^{\text{img}}$ is given as
-$\frac{S^{\text{real}}\cdot f_{1}}{D}=S^{\text{img}}_{1}\neq S^{\text{img}}_{2}=\frac{S^{\text{real}}\cdot f_{2}}{D}$ (11)
-where $f_{1}$ and $f_{2}$ denote the focal length of camera capturing the same scene.
-Taken to the extreme, for the exact same object image (same size $S^{\text{img}}$), the network needs to predict arbitrarily different distances in world coordinates due to the focal length change.
-This is clearly impossible, and this phenomenon confuses the network both in training as well as at inference time.
-When combining multiple datasets that were not taken with a single camera or even a single dataset that used different cameras to capture the dataset [38], it is a pre-requisite to be able to accommodate focal length differences.
-Figure 6: Canonical Object Space.
-One scene viewed with two different cameras with different parameters.
-As the imaging size of the same object is different, 3D labels are scaled accordingly to adapt for focal length change.
-The predictions and loss computations are all calculated in the Canonical Object Space, and the final predictions are obtained after scaling back based on (known) camera parameters.
-Inspired by Metric3D [3] where depth maps are normalized into a uniform space, our method presents Canonical Object Space (COS) where3D bounding boxes are trained to be invariant to changes in the focal length by choosing a single canonical focal length and then transforming the labels (objects) into COS.
-Instead of transforming whole images or depth maps, as in Metric3D, only the $x$,$y$,$z$ coordinates of all pseudo-ground truth labels are transformed.
-To transform labels, the scaling parameter $\omega_{i}$ is calculated as
-$\omega_{i}=\frac{f^{C}}{f_{i}}$ (12)
-where $f_{i}$ stands for the focal length of the frame $i$ and $f^{C}$ is the canonical focal length. Given the scaling parameter $\omega_{i}$, the object instance $(x,y,z)$ is transformed into COS as
-$x^{C}=x\cdot\omega_{i}\qquad y^{C}=y\cdot\omega_{i}\qquad z^{C}=y\cdot\omega_{i}$ (13)
-The model then operates in the COS and is directly supervised by the transformed labels (objects).
-During inference, all predictions are transformed from COS by dividing the depth prediction by scaling parameter $\omega_{j}$ of the inference frame $j$
-$x=\frac{x^{C}}{\omega_{j}}\qquad y=\frac{y^{C}}{\omega_{j}}\qquad z=\frac{z^{C}}{\omega_{j}}$ (14)
-This very simple transformation enables the network to train on multiple datasets where focal length differs and also allows the network to work in unseen camera setup (focal length) because $\omega_{j}$ is calculated at inference time (see table III – 0% fine-tuning row).
-Note that in the training process, it is necessary to adapt for data augmentations that affect the perceived focal length of the image, such as image scaling, and for the fact that the detector resizes the image into a constant input size by adjusting the perceived focal length of the image subject to the applied augmentations and the resizing.
-IV Experiments
-Datasets
-We use two public datasets for the evaluation – KITTI [8] and KITTI-360 [39].
-KITTI dataset setup has four front-facing cameras, IMU and LiDAR.
-KITTI-360 setup uses two front-facing cameras and two fish-eye cameras facing sides, IMU and LiDAR.
-The cameras and their parameters differ in both datasets.
-For the KITTI dataset, the same training (3712 samples) and validation (3769 samples) splits as [2, 40, 29, 31] are used.
-For the KITTI-360 dataset, the training (6 sequences, 44178 frames), validation (2 sequences, 1115 frames) and testing (1 sequence, 2459 frames) splits from VSRD [4] are used.
-It is worth noting that, unlike VSRD, we are able to use all frames of the training set as our method is much faster.
-Implementation
-We used off-the-shelf MViT2-Huge [41] in the Detectron2 framework [42] trained on MS-COCO [43] as the 2D object detector and Metric3DV2-giant [9] for metric depth estimation.
-We use our pseudo-labels to train MonoDETR [19] monocular 3D object detector, using AdamW [44] optimiser with learning rate and weight decay equal to 0.0002 and 0.0001, respectively, while keeping other hyper-parameters as in  [19].
-The canonical focal length is equal to 500.
-Aggregation is done over 100 frames, the $p$-score threshold for stationary/moving classification is 0.0001, and the minimum threshold for $d_{real}$ is 5 meters.
-The steepness parameter $\alpha$ in the Saturated Closeness Criterion is 10.
-IV-A Results
+Figure 5: Generated 3D labels are transformed into Canonical Object Space to accommodate for different focal lengths; the resulting model is therefore trained to output canonical coordinates.
+At inference time, model predictions are transformed back to world coordinates, using (known) camera parameters.
 Training supervision 3D Object detection accuracy Labelling
@@
 Methods without human annotations
-Ours 38.41/29.98 35.26/27.56 50.84/42.72 49.22/46.59 5 sec
-TABLE I: Evaluation results of monocular 3D object (vehicle) detection on the KITTI-360 test set.
+MonoSOWA (Ours) 38.41/29.98 35.26/27.56 50.84/42.72 49.22/46.59 1311 ms
+TABLE I: Evaluation results of monocular 3D object (vehicle) detection on the KITTI-360 [42] test set.
 * denotes that given data modality was not used by the method itself, but it was required in the annotation process to create human 3D box annotations.
+III-D Canonical Object Space (COS)
+Camera focal length is a crucial parameter for precise 3D position estimation, but monocular 3D detectors have no direct way to compensate for changing focal length because the same object at the same distance from the camera will have different size in the image, depending on the focal length.
+Taken to the extreme, for the exact same object image measured in pixels, the network needs to predict arbitrarily different distances in world coordinates, depending on camera focal length.
+This phenomenon confuses the network during training, and it is therefore necessary to solve this problem in order to be able to train a network on data captured by different cameras.
+Inspired by methods such as Metric3D [5] and Omni3D [6], our auto-labelling method uses Canonical Object Space (COS) where 3D bounding boxes are trained to be invariant to changes in the focal length by choosing a single canonical focal length and then transforming the generated labels (objects) into COS.
+Instead of transforming whole images or depth maps as in Metric3D, only the $x,y,z$ coordinates of pseudo-ground truth labels are transformed directly when creating the pseudo-labels.
+To transform labels, the scaling parameter $\omega_{i}$ is calculated as
+$\omega_{i}=\frac{f^{C}}{f_{i}},$ (8)
+where $f_{i}$ stands for the focal length of the frame $i$ and $f^{C}$ is the canonical focal length. Given the scaling parameter $\omega_{i}$, the object instance $(x,y,z)$ is transformed into COS as
+$x^{C}=x\cdot\omega_{i}\qquad y^{C}=y\cdot\omega_{i}\qquad z^{C}=y\cdot\omega_{i}.$ (9)
+The model then operates in the COS and is directly supervised by the transformed labels (objects).
+During inference, all predictions are transformed from COS using the scaling parameter $\omega_{j}$ of the inference frame $j$ (see fig.
+5).
+This very simple transformation enables the network to train on multiple datasets where focal length differs and also allows the network to work in unseen camera setup (focal length) because $\omega_{j}$ is calculated during inference.
+Note that in the training process, it is necessary to adapt for data augmentations that affect the perceived focal length of the image, such as image scaling, and for the fact that the detector resizes the image into a constant input size by adjusting the perceived focal length of the image subject to the applied augmentations and the resizing.
+IV Experiments
+APBEV/AP3D@0.5
+Method Published Easy Hard
+Autolabels [2] ICCV2021 51.85/4.65 46.10/2.92
+VSRD [4] CVPR2024 47.12/35.25 43.91/32.64
+Ours - 61.17/47.07 51.92/45.51
+TABLE II: Auto-labelling AP on KITTI-360 [42] training set.
+Training supervision 3D Object detection accuracy (Level 2)
+Human Labels APBEV@0.5 AP3D@0.5
+Method LiDAR Masks 3D boxes All 0m-30m 30m-50m 50m-$\infty$ All 0m-30m 30-50m 50m-$\infty$
+DEVIANT [43] yes* yes 10.29 26.75 4.95 0.16
+MonoDETR [20] yes* yes 23.63 39.10 20.95 4.70 21.41 37.89 18.61 3.69
+MonoSOWA (Ours) 18.98 33.51 12.02 4.55 13.46 24.65 5.87 4.55
+TABLE III: Monocular 3D object (vehicle) detection AP on Waymo [44] validation set.
+Datasets.
+We use three public datasets for evaluation – KITTI [10], KITTI-360 [42] and Waymo [44].
+In the KITTI dataset, we employ the right RGB camera; in the KITTI-360 dataset, we use the right perspective camera; and in Waymo, we utilize the front-facing camera.
+Note that the camera parameters – their focal length, resolution, and FOV – are different in each dataset.
+In KIITI, we use the same training (3,712 samples) and validation (3,769 samples) splits as [2, 45, 29, 30] are used.
+In KITTI-360 dataset, training (6 sequences, 44,178 frames) and testing (1 sequence, 2,459 frames) splits from VSRD [4] are used.
+Waymo experiments follow the official training (798 sequences, 158,080 frames) and validation (202 sequences, 39,988 frames) splits.
+Implementation.
+We used off-the-shelf MViT2-Huge [40] in the Detectron2 framework [46] trained on MS-COCO [47] as the 2D object detector and Metric3DV2-giant [11] for metric depth estimation.
+We use our pseudo-labels to train MonoDETR [20] monocular 3D object detector, using AdamW [48] optimiser with learning rate and weight decay equal to 0.0002 and 0.0001, respectively, while keeping other hyper-parameters as in  [20].
+The canonical focal length $f_{c}$ is equal to 750.
+Aggregation is done over 100 frames, the $T_{z}$ threshold for stationary/moving classification is 0.2, and the minimum threshold $T_{m}$ is 5 meters.
+The steepness parameter $\alpha$ in the Saturated Closeness Criterion is 10 (see Supplementary material for ablations).
+Note that we used the same hyper-parameter settings for all three datasets.
+IV-A Results
+Training Fine-tuning APBEV/AP3D@0.7
+pseudo labels human labels Easy Moderate Hard
+VSRD [4] 0% 0.002/0.001 0.004/0.001 0.005/0.002
+Ours 24.46/8.24 20.03/6.84 16.14/6.36
+VSRD 25% 31.72/21.76 22.32/15.43 18.86/12.55
+Ours 39.99/32.64 30.92/25.06 26.47/20.35
+VSRD 50% 43.44/31.05 31.54/21.48 27.17/17.93
+Ours 44.32/33.70 32.97/25.65 27.52/23.08
+VSRD 75% 42.58/32.95 31.08/24.68 27.19/21.38
+Ours 43.32/34.11 32.48/25.68 27.39/23.05
+- 100% 37.99/29.36 26.76/20.64 23.02/17.30
+TABLE IV: Monocular 3D object detection evaluated on KITTI [10] validation set. Pre-training on KITTI-360 [42] pseudo-labels and fine-tuning on a fraction of the KITTI training set.
+KITTI-360.
+In the first experiment, we generate pseudo-labels on the KITTI-360 [42] dataset and train MonoDETR [20] with these pseudo-labels.
+We then compare our trained model with previously reported results of other trained monocular 3D object detectors.
+As shown in table I, our method outperforms current state-of-the-art weakly-supervised method VSRD [4] both in Bird’s eye view (BEV) and 3D average precision at 0.5 IoU, despite the fact our method does not use any human labels.
+In the 0.3 IoU metric, our method outperforms both WeakM3D [1] and Autolabels [2] by a significant margin both for BEV and 3D and it achieves competitive results with VSRD [4].
+The remaining gap in the IoU 0.3 metric is caused by KITTI-360 human labels which are amodal (they include occluded or invisible object parts), which significantly helps VSRD that uses them in its training.
+Our method instead relies on a 2D detector, which only outputs boxes based on what is actually visible in the image, but as shown in table X, when identical inputs are used, our method significantly outperforms VSRD in all metrics.
+Note that our auto-labelling pipeline takes, on average, 1.3 seconds per frame (including all the models it uses as inputs - see table IX) – a 700x speed-up.
+We also directly compare all generated labels with human labels on the KITTI-360 training set (see table II).
+We show significantly improved accuracy of our pseudo-labels compared to previous auto-labelling methods [4, 2], both in 3D and BEV.
+Waymo.
+In the second experiment, we evaluate our method on Waymo dataset [44], where we again train MonoDETR [20] detector using only our pseudo-labels generated for Waymo training set and evaluate on Waymo validation set (see table III).
+Despite using no human annotations, our method achieves competitive performance to methods using human annotations.
+Note that we are unable to directly compare to VSRD [4] on this dataset, because the training set has 150k frames, and it is therefore practically infeasible to auto-label the data using VSRD (15 minutes per frame) in any reasonable time frame.
+Pseudo-labels for pre-training.
+In the following experiments, the labels generated by our auto-labelling pipeline are used as a pre-training step and then only a small fraction of human labels is used to further fine-tune the model.
+Training Fine-tuning APBEV AP3D@0.5
+pseudo labels human labels
+Ours 1% 21.59 18.43
+- 1% 15.55 13.81
+Ours 5% 22.56 20.68
+- 5% 21.08 18.12
+Ours 10% 22.80 20.65
+- 10% 21.69 18.53
+Ours 15% 23.20 20.88
+- 15% 22.63 20.60
+- 100% 23.63 21.41
+TABLE V: Monocular 3D object detection (vehicle - Level 2) on Waymo validation set. Training on Waymo training set using pseudo-labels and fine-tuning on a fraction of human labels.
+First, in table IV we show that pre-training on a larger dataset (KITTI-360) auto-labelled by our method and then fine-tuning on the target domain (KITTI) significantly improves the average precision when compared to simply training on the target domain from scratch.
+With only 25% of human labels, our method outperforms traditional fully-supervised detector trained from scratch with 100% of human labels, and the gap grows as more human labels are used.
+Also note that our method significantly outperforms current state-of-the-art VSRD [4] by a large margin.
+Second, we show on a large-scale dataset (Waymo) that pre-training with our method can save human annotations costs.
+In this scenario, we assume the annotation budget is limited and therefore it’s not possible to label all images in the training set.
+We show, that when only a fraction of images are labelled by a human – e.g.
+only 5%, using our pseudo-labels as a pre-training step leads to consistent improvements in detection accuracy at no annotation cost, compared to only using available human labels (see table V).
+Also, when using only 15% of human labels and our method as a pre-training step, the accuracy is almost identical to using all human labels; this can also be viewed that using our method leads to saving 85% of annotation costs for a very small drop in accuracy.
 Dataset Labels APBEV/AP3D@0.5 APBEV/AP3D@0.3
 for training Easy Moderate Hard Easy Moderate Hard
-KITTI pseudo 59.76/51.55 44.08/37.09 36.99/33.15 73.38/72.70 57.23/56.30 48.59/47.70
+KITTI pseudo 61.24/53.22 45.66/40.88 38.67/34.68 77.69/76.07 63.39/58.20 55.70/54.46
 K360 pseudo 57.62/47.39 43.86/38.33 37.05/32.23 75.83/74.40 63.69/61.56 55.75/53.85
-KITTI + K360 pseudo 62.28/53.04 46.38/40.79 42.38/35.10 79.70/78.48 65.85/63.97 57.41/55.46
+KITTI + K360 pseudo 64.29/58.97 50.78/44.23 44.56/40.58 78.27/77.71 65.86/64.85 62.35/56.67
 KITTI human 67.44/65.09 53.46/47.39 46.80/44.58 80.30/79.72 67.16/65.87 59.54/58.83
-TABLE II: Monocular 3D object detection accuracy on the KITTI validation set, depending on which datasets/labels are used in training.
-KITTI-360.
-The first evaluation and comparison with prior work is on KITTI-360 [39] as shown in table I.
-Our method shows outstanding results on both the Bird’s eye view (BEV) and 3D at 0.5 IoU as it outperforms current state-of-the-art weakly-supervised method VSRD [4], despite the fact our method does not use any human labels.
-In the 0.3 IoU metric both for BEV and 3D, our method outperforms both WeakM3D [1] and Autolabels [2] by a significant margin and it achieves a competitive AP as VSRD [4].
-We believe that this is likely caused by the detection failures, class ambiguities or that KITTI 2D bounding boxes are amodal as using consistent ground-truth 2D labels significantly helps VSRD.
-Cross-dataset evaluation
-Thanks to the proposed Canonical Object Space, it is possible to train on multiple merged datasets.
-We, therefore, combine both KITTI and KITTI-360 datasets training subsets, generate our pseudo-labels, train a single model and evaluate it on the KITTI validation subset (see table II).
-We show that generating pseudo-labels for two different datasets and then performing training on the joint significantly improves average precision both for 0.3 and 0.5 IoU on BEV and 3D.
-The gain comes from the additional amount of data present in the KITTI-360 dataset.
-We also demonstrate the ability of our method to work out-of-the-box in previously unseen camera setup (see table II - row K360), where the model was trained using pseudo-labels on KITTI-360 but evaluated on in this case unseen KITTI dataset.
-We note that considering IoU 0.3, the accuracy of a model trained on KITTI-360 is actually even slightly higher than the model trained on the original domain (KITTI), which we contribute simply to the fact that KITTI-360 contains more training data.
-Semi-supervised
-GT APBEV/AP3D@0.7
-Method Ratio Easy Moderate Hard
-VSRD [4] 0% 0.002/0.001 0.004/0.001 0.005/0.002
-Ours 24.46/8.24 20.03/6.84 16.14/6.36
-VSRD [4] 25% 31.72/21.76 22.32/15.43 18.86/12.55
-Ours 39.99/32.64 30.92/25.06 26.47/20.35
-VSRD [4] 50% 43.44/31.05 31.54/21.48 27.17/17.93
-Ours 44.32/33.70 32.97/25.65 27.52/23.08
-VSRD [4] 75% 42.58/32.95 31.08/24.68 27.19/21.38
-Ours 43.32/34.11 32.48/25.68 27.39/23.05
-MonoDETR [19] 100% 37.99/29.36 26.76/20.64 23.02/17.30
-TABLE III: 3D object detection accuracy on KITTI validation set depending on the fraction of human labels used to fine-tune the model.
-In this experiment, MonoDETR is trained jointly on KITTI and KITTI-360, while all labels are generated by our auto-labelling pipeline, as a pre-training step.
-Next, only a given fraction of human labels is used to fine-tune the model.
-Firstly, as seen in table III, thanks to COS our method works out-of-the-box without any fine-tuning (0% human labels), unlike previous methods, showing there is no need to fine-tune using human labels while moving to different domains.
-Secondly, pre-training on a larger dataset auto-labelled by our method and then fine-tuning on the target domain significantly improves the average precision over simply training on the target domain from scratch.
-With only 25% of the ground truth labels, our method outperforms traditional fully-supervised detector trained from scratch with 100% of human labels, and the gap grows as more human labels are used.
-Lastly, our method significantly outperforms current state-of-the-art VSRD [4] by a large margin, showing that pre-training on larger datasets is a promising direction for further research.
-Figure 7: Example 3D car detections of our model trained without using human annotations on both KITTI [8] and KITTI-360 [39].
+TABLE VI: MonoDETR [20] AP on the KITTI [10] validation set, depending on which datasets/pseudo-labels are used in training.
+Cross-dataset evaluation.
+Thanks to using Canonical Object Space, it is possible to combine datasets with different camera setups.
+First, we demonstrate that unlike previous work [4], our method works out-of-the-box in previously unseen camera setup (see table VI - row K360 and table IV - row 0%), where the model was trained using pseudo-labels on KITTI-360 but evaluated on in this case unseen KITTI dataset.
+We also combine both KITTI and KITTI-360 datasets training subsets, generate our pseudo-labels, train a single model and evaluate it on the KITTI validation subset (table VI – row KITTI+K360).
+Combining pseudo-labels from KITTI+K360 achieves comparable accuracy to using human labels of KITTI, and even surpasses the accuracy of training with human labels in the Hard category in the 0.3 metric.
+Figure 6: Example of 3D car detections by MonoDETR [20] trained without human annotations on KITTI [10] (top), KITTI-360 [42] (middle) and Waymo [44] (bottom)
+Qualitative analysis.
+We provide qualitative analysis in fig. 6 and in the Supplementary material in fig. 7 and fig. 8.
 IV-B Ablations
+In this section, we present two ablation experiments we deemed the most important, and, due to lack of space, we kindly refer the reader to the Supplementary material for further ablations.
+The auto-labelling pipeline ablations use KITTI [10] training set for evaluation as it best measures the quality of the auto-generated labels, while the ablations involving a trained model (MonoDETR) use the respective KITTI training/validation sets for training and evaluation.
+First, we inspect the impact of the proposed Local Object Motion Model and Saturated Closeness Criterion.
+As seen in  table VII, both components are crucial to achieve the best accuracy of generated pseudo-labels.
+LOMM SCC APBEV/AP3D@0.5
+Easy Moderate Hard
+20.41/17.40 18.22/15.88 19.99/13.75
+20.37/17.16 18.26/15.76 20.04/13.66
+35.89/29.05 26.32/17.43 28.87/19.19
+39.22/32.29 33.00/23.14 31.09/21.34
+TABLE VII: Local Object Motion Model (LOMM) and Saturated Closeness Criterion (SCC) ablation on the KITTI [10] training set.
+In the second ablation, we don’t use Canonical Object Space and instead naively combine KITTI and KITTI-360 datasets.
+As expected, the average precision decreases when KITTI-360 data are appended to KITTI (see table VIII), because the network focuses on camera setup of the larger K360 dataset, and it is only with COS enabled that the network actually benefits from the added data.
+Note that only in this ablation, we use human labels to eliminate any effects of pseudo-labels.
 Dataset APBEV/AP3D@0.3
@@
 KITTI+K360 82.07/81.25 67.14/66.16 62.47/61.02
-TABLE IV: Canonical Object Space (COS) ablation on KITTI validation set under full supervision from human labels.
-Canonical Object Space
-As seen in table IV, disabling COS and training on both datasets jointly decreases the average precision of both BEV and 3D.
-It may be surprising that when training on both KITTI+K360, the model is not broken completely – we speculate that the network is able to learn to classify from which of the two datasets the image comes from and adjusts its outputs accordingly; such an approach obviously does not generalize well.
-Note that only in this ablation, we use human labels to eliminate any effects of pseudo-labels.
-Number of frames used in aggregation
-No. of APBEV/AP3D@0.5
-frames Easy Moderate Hard
-$\pm$ 1 53.26/49.35 41.34/35.86 34.77/29.14
-$\pm$ 10 59.76/61.55 44.08/37.09 36.99/33.15
-$\pm$ 30 59.22/51.18 42.65/36.64 35.91/29.77
-TABLE V: Number of frames used in the aggregation process ablation on KITTI validation set.
-table V shows the ablation study on the number of frames used in the aggregation process.
-Perhaps surprisingly, aggregating over longer sequences does not yield better performance.
-The aggregation can recover from the error caused by Metric3D [3, 9] as in fig.
-4.
-We speculate that the distant car poses a significant challenge for the detector itself, therefore creating precise 3D labels for such objects does not translate to better 3D detection accuracy.
-APBEV/AP3D@0.5
-$\mathbf{\alpha}$ Easy Moderate Hard
-1 53.02/48.09 41.29/34.92 34.48/28.66
-5 55.14/50.83 42.46/36.59 35.92/29.85
-10 56.56/48.90 41.32/35.00 34.53/28.76
-15 55.78/51.86 42.48/36.63 35.72/29.81
-TABLE VI: Steepness parameter $\alpha$ of the Saturated Closeness Criterion ablation on KITTI validation set.
-Steepness parameter in Saturated Closeness Criterion
-In this ablation, we explore possible values for the steepness parameter $\alpha$ in the Saturated Closeness Criterion, which controls what pseudo-LIDAR points are discarded as outliers.
-As shown in table VI, $\alpha$ equal to 10 achieves the best performance.
+TABLE VIII: Canonical Object Space (COS) ablation using human labels, evaluated on KITTI validation set.
 V Conclusion
@@
 The method exploits temporal consistency in video sequences to automatically create 3D labels of objects (cars) and, as such, does not require human labels or additional sensors such as LiDAR.
-The method is able to compensate for camera focal length differences to significantly improve its scalability across multiple datasets, and furthermore, it works out of the box in previously unseen camera setups.
-The method is evaluated on two large-scale public datasets, where despite using no human labels, it outperforms prior work by a significant margin both for BEV and 3D when using the stricter 0.5 IoU evaluation.
-Last but not least, the method shows its versatility by also being a powerful pre-training method for fully-supervised monocular 3D detection, where the final accuracy increases by approximately 6 percent points over traditional object detector.
+The method uses newly proposed Local Object Motion Model (LOMM) to disentangle object movement source between subsequent frames, is approximately 700 times faster than current state of the art and is able to compensate for camera focal length differences to improve scalability across multiple datasets.
+The method is evaluated on three large-scale public datasets, where despite using no human labels, it outperforms prior work by a significant margin both for BEV and 3D when using the stricter 0.5 IoU evaluation.
+The method also shows its versatility by being a powerful pre-training tool for fully-supervised monocular 3D detection, and last but not least, shows that combining pseudo-labels from multiple datasets can achieve comparable accuracy to using human labels from a single dataset.
 The main limitation is detecting objects far away, which is not primarily due to the auto-labelling procedure but due to the inherent limitation of the trained monocular detector to infer the distance of objects which are only a couple of pixels high.